OpenStudy (anonymous):

Someone answer 5ft=_____in math please solve it

OpenStudy (anonymous):

Inches

OpenStudy (anonymous):

thats in

OpenStudy (jamesj):

1 foot = 12 inches 2 feet = 2*12 inches = 24 inches 3 feet = 3*12 inches = 36 inches Hence 5 feet = what?

OpenStudy (anonymous):

how do you put it in inches all math

OpenStudy (jamesj):

2 feet = 2*12 inches = 24 inches 3 feet = 3*12 inches = 36 inches 4 feet = 4*12 inches = 48 inches Hence 5 feet = what?

OpenStudy (anonymous):

how many inches are in 5 feet its customary units of legth you know??!!1

OpenStudy (anonymous):



OpenStudy (jamesj):

Look at how you do it. I am showing you 2 feet = 2x12 inches = 24 inches 3 feet = 3x12 inches = 36 inches 4 feet = 4x12 inches = 48 inches Hence 5 feet = 5x12 inches = ___ inches

OpenStudy (anonymous):

60

OpenStudy (jamesj):

yes

OpenStudy (anonymous):

How bout 3mi. = ____?____ ft.

OpenStudy (jamesj):

How many feet in 1 mile?

OpenStudy (anonymous):

5,280 ft.

OpenStudy (jamesj):

1 mile = 5,280 ft 2 miles = 2x5,280 ft = 10,560 ft 3 miles = ...

OpenStudy (anonymous):

by 3 is 15660

OpenStudy (jamesj):

Not quite. 3 miles = 3x5,280 ft = ...... ft

OpenStudy (anonymous):

Mt answer changes 15840

OpenStudy (anonymous):

my*

OpenStudy (jamesj):

Yes. 3 miles = 3x5,280 ft = 15,840 ft

OpenStudy (anonymous):

HOw bout 24ft=___yd

OpenStudy (jamesj):

How many feet in a yard?

OpenStudy (anonymous):

3

OpenStudy (jamesj):

Right. So you now figure out the answer. 1 yard = 3 feet 2 yards = 6 feet 3 yards = 9 feet 4 yards = 12 feet ...

OpenStudy (anonymous):